For the United States, the mean monthly Internet bill is $32.79 per household (CNBC, January 18, 2006). A sample of 50 households in a southern state showed a sample mean of $30.63. Use a population standard deviation of $5.60. Formulate hypothesis for a test to determine whether the sample data support the conclusion that the mean monthly internet bill in the southern state is less than the national mean $32.79.


1
Expert's answer
2021-03-08T19:19:45-0500

"n = 50 \\

\bar{x} = 30.63 \\

σ = 5.60 \\

α = 0.01 \\

H_0: \; μ ≥ 32.79 \\

H_1: \; μ< 32.79 \\

Z = \frac{\bar{x}-μ}{σ/ \sqrt{n}} = \frac{30.63-32.79}{5.60/ \sqrt{50}} = -2.73 \\

P = P(Z<-2.73) = 0.0032 \\

P < 0.01 \\

Reject \; H_0."